x-intercepts for the parabola defined by the equation below.

y = 3x2 + 18x + 15

Answer by stanbon(75887)

x-intercepts for the parabola defined by the equation below.
y = 3x^2 + 18x + 15
-------------------------------------------
Solve: 3x^2 + 18x + 15 = 0
3x^2 + 3x + 15x + 15 = 0
3x(x+1) + 15(x+1) = 0
(x+1)(3x+15) = 0
x = -1 and x = -5 are the x-intercepts
==========================================
Cheers,
Stan H.
=========